# LoRA-Mixer: Coordinate Modular LoRA Experts Through Serial Attention Routing

### v1

@@ 5 Conclusion and Discussion
This paper introduces LoRA-Mixer, a flexible and architecture-agnostic MoE framework for combining LoRAs, adapting LLM for multitask. It improves the performance of Transformer and SSM models by replacing the projection layer with a dynamically routed LoRA experts. Through a two-stage training paradigm, LoRA-Mixer decouples expert learning from routing, enabling specialization and task awareness. To address the problem of overly uniform auxiliary losses, we propose RSL, which balances expert load while improving routing selectivity. The framework enables efficient router training with minimal data and supports cross-domain reuse of LoRA modules. Although LoRA-Mixer is effective, its fixed top-$K$ routing may limit adaptability to ambiguous inputs. Uniformly applying it across all layers can also introduce redundancy, as different layers capture different information. Future work will explore dynamic or differentiable routing and adaptive integration to apply LoRA-Mixer only where most beneficial.
Appendix A Experiment Result
@@
As shown in Table LABEL:r16,Table 10 and Table  11, within a certain range, as $r$ increases, the performance of the model can be improved to a certain extent. Our method is not only better than the basic model, but even better than the fine-tuned basic model on most tasks. This shows that our method can effectively combine existing knowledge through dynamic expert combination to form a more "intelligent" model.
Appendix B Falcon-Mamba Architecture Analysis
$h^{\prime}(t)$ $=\mathbf{A}h(t)+\mathbf{B}x(t),$ (9)
$y(t)$ $=\mathbf{C}h(t).$
$\mathbf{\overline{A}}$ $=\exp\left(\mathbf{\Delta A}\right),$ (10)
The discrete state-space equation with a step size of $\Delta$ becomes:
$h_{t}$ $=\mathbf{\overline{A}}h_{t-1}+\mathbf{\overline{B}}x_{t},$ (11)
$y_{t}$ $=\mathbf{C}h_{t}.$
$y$ $=x\otimes\mathbf{\overline{K}}.$
$x\xrightarrow{x\text{-proj}}(\Delta,B,C)$
$\Delta\xrightarrow{dt\text{-proj}}\Delta_{discretized}$
These discretized values—$\Delta_{discretized}$, $A$, $B$, $C$, and $D$—are then fed into the selective scanning module for processing. This architectural design of Falcon-Mamba fully enables the application of LoRA-Mixer specifically tailored for the projection layer. For a comprehensive understanding of the training process, please refer to.
Appendix C Theoretical Justification of RSL Loss
We provide a theoretical analysis of the proposed RSL loss and contrast it with the conventional auxiliary loss. Our goal is to demonstrate that RSL naturally promotes input-aware, expert-specialized routing with improved data efficiency.
C.1 Preliminaries
Let the router output a softmax distribution $p(x)=[p_{1}(x),\dots,p_{K}(x)]\in\Delta^{K-1}$ over $K$ experts for a token $x$, where
$p_{i}(x)=\frac{\exp(G_{i}(x))}{\sum_{j=1}^{K}\exp(G_{j}(x))},\quad\sum_{i=1}^{K}p_{i}(x)=1.$ (13)
We define the expected routing probability and top-1 selection frequency as:
$\bar{p}_{i}$ $=\mathbb{E}_{x\sim\mathcal{D}}[p_{i}(x)],$ (14)
$\bar{f}_{i}$ $=\mathbb{E}_{x\sim\mathcal{D}}\left[\mathbb{I}(i=\arg\max_{j}p_{j}(x))\right],$ (15)
where $\bar{p}_{i}$ represents the average routing intention, and $\bar{f}_{i}$ represents the empirical usage under hard top-1 routing.
C.2 Auxiliary Loss and Its Implicit Bias
The standard auxiliary loss encourages load balancing by aligning the average routing with actual expert usage:
$\mathcal{L}_{\text{aux}}=\alpha\sum_{i=1}^{K}\bar{p}_{i}\cdot\bar{f}_{i}.$ (16)
Proposition 1 (Equilibrium of Auxiliary Loss).
Under the constraint $\sum_{i=1}^{K}\bar{p}_{i}=1$, the minimum of $\mathcal{L}_{\text{aux}}$ is attained when $\bar{p}_{i}=\bar{f}_{i}$ for all $i$.
Proof. Using the Lagrangian method:
$\mathcal{L}=\sum_{i=1}^{K}\bar{p}_{i}\bar{f}_{i}-\lambda\left(\sum_{i=1}^{K}\bar{p}_{i}-1\right)$
Taking partial derivatives:
$\frac{\partial\mathcal{L}}{\partial\bar{p}_{i}}=\bar{f}_{i}-\lambda=0\Rightarrow\bar{f}_{i}=\lambda,\quad\forall i.$
Thus, all $\bar{f}_{i}$ are equal, implying uniform distribution: $\bar{f}_{i}=\frac{1}{K}$, hence $\bar{p}_{i}=\frac{1}{K}$.
This shows that the auxiliary loss alone biases the router toward uniform expert activation, regardless of input characteristics.
C.3 RSL Loss: Entropy-Regularized Routing
To promote more input-sensitive routing, we introduce an entropy-regularized objective:
$\mathcal{L}_{\text{RSL}}=\mathcal{L}_{\text{aux}}-\lambda\cdot\mathbb{E}_{x\sim\mathcal{D}}[\mathcal{H}(p(x))],$ (17)
where the token-level entropy is defined as:
$\mathcal{H}(p(x))=-\sum_{i=1}^{K}p_{i}(x)\log p_{i}(x).$ (18)
This term encourages the router to assign higher weights to fewer experts, effectively breaking uniformity and encouraging selective specialization.
C.4 Gradient Analysis and Token-Awareness
We derive the entropy gradient w.r.t. routing score $p_{i}(x)$:
$=-\log p_{i}(x)-1+\mu,$ (20)
where $\mu$ is the Lagrange multiplier due to the simplex constraint.
Thus, the total gradient of the RSL loss becomes:
This shows that RSL introduces token-level gradient signals via $\log p_{i}(x)$, unlike the auxiliary loss, which propagates only global gradients.
C.5 Token-Awareness via Routing Variance
To quantify input-aware routing, we define:
We say the routing is token-aware if $\text{Var}(p(x))>\epsilon$ for some $\epsilon>0$. The auxiliary loss tends to reduce this variance (driving uniform routing), while RSL encourages high variance and peaked distributions aligned with input semantics.
C.6 Conclusion
The RSL loss incorporates an entropy-based regularizer that mitigates the uniformity bias of auxiliary loss. By injecting token-level gradient signals and promoting routing variance, RSL enables input-aware, specialized, and discriminative expert assignments. This property is especially beneficial in data-scarce regimes, where each token’s contribution to routing must be maximally leveraged. Note that RSL is fully compatible with auxiliary loss; it can be viewed as a strict generalization that stabilizes training while encouraging specialization.
@@ Appendix D Balance loss visualization
Appendix E Details of LoRA modules of Flan-T5.

### v2

@@ 4.5 Ablation Study
This paper introduces RSL, designed to address the issue of excessively uniform auxiliary loss. It achieves highly selective routing while achieving overall load balancing, and enables efficient routing training with minimal data, providing strong support for the reuse of LoRAs modules. To verify the effectiveness of RSL, we propose the LoRA-Mixer framework and integrate RSL. LoRA-Mixer is a flexible and architecture-agnostic MoE framework for combining LoRAs and adapting LLMs to multitasking. Unlike other methods that indiscriminately insert MoEs or completely replace attention or FFN modules, LoRA-Mixer only adapts the core projection layer, improving the performance of Transformers and SSM models. Although LoRA-Mixer is effective, its fixed top-$K$ routing may limit adaptability to ambiguous inputs. Uniformly applying it across all layers can also introduce redundancy, as different layers capture different information. Future work will explore dynamic or differentiable routing and adaptive integration to apply LoRA-Mixer only where most beneficial.
6 Acknowledgments
This research was supported by the National Natural Science Foundation of China (Nos. 62272184 and 62402189), the China Postdoctoral Science Foundation (Nos. 2024M751012, 2025T180429, and GZC20230894), and the Hubei Provincial Postdoctoral Research Funding Program (No. 2024HBBHCXB014). The computational work was performed on the high-performance computing platform at Huazhong University of Science and Technology.
Appendix A Appendix
A.1 Optimization and Convergence Analysis.
We analyze the empirical counterpart of $\mathcal{L}_{\text{RSL}}$ on a dataset $S=\{x_{j}\}_{j=1}^{n}$:
$\widehat{\mathcal{L}}_{\text{RSL}}(S;\{p(x_{j})\})$ $=\;\alpha\sum_{i=1}^{K}\widehat{\bar{p}}_{i}\cdot\widehat{\bar{f}}_{i}\;-\;\frac{\lambda}{n}\sum_{j=1}^{n}\mathcal{H}\!\big(p(x_{j})\big),$
$\widehat{\bar{p}}_{i}$ $=\frac{1}{n}\sum_{j=1}^{n}p_{i}(x_{j}),\qquad\widehat{\bar{f}}_{i}=\frac{1}{n}\sum_{j=1}^{n}\mathbb{I}\!\big(i=\arg\max_{\ell}p_{\ell}(x_{j})\big).$ (13)
Smooth surrogate for optimization.
For optimization analysis we replace the hard top-$1$ frequency by a standard smooth surrogate:
Assumption 1 (Smoothed usage surrogate).
Replace $\widehat{\bar{f}}_{i}$ in equation 13 by $\widehat{\bar{s}}_{i}$, where $s_{i}^{\tau}(x)=\mathrm{softmax}(G(x)/\tau)_{i}$ for some $\tau>0$ and $\widehat{\bar{s}}_{i}=\frac{1}{n}\sum_{j}s_{i}^{\tau}(x_{j})$. We assume the composite $\sum_{i}\widehat{\bar{p}}_{i}\,\widehat{\bar{s}}_{i}$ is convex and $L$-smooth in $\{p(x_{j})\}$ on the product simplex $\Delta^{n}$. [^1] This includes the commonly used choice $\widehat{\bar{s}}_{i}=\widehat{\bar{p}}_{i}$, giving the auxiliary term $\alpha\sum_{i}\widehat{\bar{p}}_{i}^{2}$, which is convex since each $\widehat{\bar{p}}_{i}$ is linear in $\{p(x_{j})\}$.
Negative-entropy–regularized objective.
Define
$F_{S}(\{p(x_{j})\})$ $:=\alpha\sum_{i=1}^{K}\widehat{\bar{p}}_{i}\cdot\widehat{\bar{s}}_{i}+\frac{\lambda}{n}\sum_{j=1}^{n}\sum_{i=1}^{K}p_{i}(x_{j})\log p_{i}(x_{j}).$ (14)
Since $\sum_{i}p_{i}\log p_{i}=\mathcal{H}(p)$, minimizing $F_{S}$ is identical to minimizing $\widehat{\mathcal{L}}_{\text{RSL}}$ with $\widehat{\bar{f}}_{i}$ replaced by $\widehat{\bar{s}}_{i}$; a change of logarithm base is absorbed into $\lambda$. We optimize $F_{S}$ over $\{p(x_{j})\}\in\Delta^{n}$ using entropic mirror descent (exponentiated gradient)
$p(x_{j})^{t+1}~\propto~p(x_{j})^{t}\odot\exp\!\Big(-\eta_{t}\nabla_{p(x_{j})}F_{S}(\{p(x_{\ell})^{t}\})\Big),\quad\text{then normalize onto }\Delta.$ (15)
Lemma 1 (Strong convexity).
The function $h(p)=\sum_{i=1}^{K}p_{i}\log p_{i}$ is $1$-strongly convex on $\Delta$ w.r.t. the $\ell_{1}$ norm. Hence the regularizer $\frac{\lambda}{n}\sum_{j=1}^{n}h\big(p(x_{j})\big)$ is $\lambda$-strongly convex on the product simplex $\Delta^{n}$. Therefore, $F_{S}$ in equation 14 is $\lambda$-strongly convex, since the usage surrogate term is convex by Assumption 1.
Lemma 2 (Gradient bounds and interior smoothness).
Under Assumption 1 there exists $G>0$ such that for all feasible $\{p(x_{j})\}$,
$\Big\lVert\nabla_{p(x_{j})}\Big(\alpha\sum_{i=1}^{K}\widehat{\bar{p}}_{i}\cdot\widehat{\bar{s}}_{i}\Big)\Big\rVert\leq G.$
Moreover,
$\Big\lVert\nabla_{p(x_{j})}\frac{\lambda}{n}\sum_{i=1}^{K}p_{i}(x_{j})\log p_{i}(x_{j})\Big\rVert\leq\frac{\lambda}{n}\,\big\|1+\log p(x_{j})\big\|,$
so if we maintain $p_{i}(x_{j})\geq\epsilon$ during training (e.g., via temperature/clipping), the right-hand side is uniformly bounded by a constant $C(\lambda,\epsilon,K)$. On the interior $\{p:p_{i}\geq\epsilon\}$, the entropy term is $L_{\text{ent}}\!\leq\!\frac{\lambda}{n\epsilon}$-smooth (w.r.t. $\ell_{2}$), hence $F_{S}$ is $L$-smooth with $L\leq L_{\text{ent}}+L_{\text{aux}}$ for some finite $L_{\text{aux}}=O(\alpha/n)$.
Theorem 1 (Convergence of EG/EMD for $F_{S}$).
Under Assumption 1 and Lemmas 12, $F_{S}$ is $\lambda$-strongly convex on $\Delta^{n}$. (a) Sublinear rate. With steps $\eta_{t}=\frac{1}{\lambda t}$, the averaged iterate $\bar{P}^{\,T}=\frac{1}{T}\sum_{t=1}^{T}P^{t}$ satisfies
$F_{S}(\bar{P}^{\,T})-F_{S}(P^{\star})\;\leq\;\mathcal{O}\!\Big(\frac{G_{\mathrm{eff}}^{2}}{\lambda\,T}\Big),$
where $G_{\mathrm{eff}}$ aggregates the uniform gradient bounds in Lemma 2. (b) Linear rate on the interior. If, in addition, training maintains $p_{i}(x_{j})\geq\epsilon$ (hence $F_{S}$ is $L$-smooth on the interior), then with any constant step $\eta\leq 1/L$ the EG iterates enjoy geometric contraction in KL divergence:
$D_{\mathrm{KL}}(P^{\star}\!\parallel P^{t+1})\;\leq\;(1-\eta\lambda)\,D_{\mathrm{KL}}(P^{\star}\!\parallel P^{t}),$
and consequently $F_{S}(P^{t})-F_{S}(P^{\star})$ also decays geometrically.
The negative-entropy term supplies curvature in the routing simplex (Lemma 1), turning the merely convex surrogate into a strongly convex objective. This stabilizes and accelerates routing optimization while preserving input-awareness through the token-level $\log p_{i}(x)$ signals (cf. the gradient form in the main RSL, Eq. equation 9).
A.2 Generalization Bound via Algorithmic Stability.
We analyze generalization under the same smoothed usage surrogate as in optimization (Assumption 1), i.e., replacing $\widehat{\bar{f}}_{i}$ by $\widehat{\bar{s}}_{i}$ (in particular, the commonly used choice $\widehat{\bar{s}}_{i}=\widehat{\bar{p}}_{i}$ yields a convex quadratic auxiliary term). We also keep the empirical objective in sample-average form so that the regularization strength does not scale with $n$.
Let $S=\{x_{j}\}_{j=1}^{n}$ and $S^{(i)}$ be $S$ with the $i$-th example replaced. Denote by $\hat{P}(S)$ a (unique) minimizer of the empirical objective
$F_{S}(\{p(x_{j})\})~=~\alpha\sum_{i=1}^{K}\widehat{\bar{p}}_{i}\cdot\widehat{\bar{s}}_{i}~+~\frac{\lambda}{n}\sum_{j=1}^{n}\sum_{i=1}^{K}p_{i}(x_{j})\log p_{i}(x_{j}),$
and similarly $\hat{P}(S^{(i)})$ for $F_{S^{(i)}}$. For reference, the population counterpart is
$F(\{p(x)\})~=~\alpha\sum_{i=1}^{K}\bar{p}_{i}\cdot\bar{s}_{i}~+~\lambda\,\mathbb{E}_{x}\!\Big[\sum_{i=1}^{K}p_{i}(x)\log p_{i}(x)\Big],\quad\bar{p}_{i}=\mathbb{E}_{x}[p_{i}(x)],~~\bar{s}_{i}=\mathbb{E}_{x}[s_{i}^{\tau}(x)].$
Assumption 2 (Lipschitz loss & bounded differences).
(i) The map $\{p(x_{j})\}\mapsto\alpha\sum_{i=1}^{K}\widehat{\bar{p}}_{i}\cdot\widehat{\bar{s}}_{i}$ is $L$-Lipschitz on the product simplex endowed with the averaged block norm $\|P\|_{\text{avg-}1}:=\tfrac{1}{n}\sum_{j}\|p(x_{j})\|_{1}$, and replacing one sample in $S$ changes this auxiliary term by at most $\tfrac{B}{n}$ (bounded-difference property). (ii) The per-example evaluation loss $\ell(p;x)$ is $\rho$-Lipschitz in $p$ (under $\|\cdot\|_{1}$ on a single simplex) and bounded in $[0,M]$.
Lemma 3 (Strong convexity under the averaged product norm).
Negative entropy $h(p)=\sum_{i=1}^{K}p_{i}\log p_{i}$ is $1$-strongly convex on $\Delta$ w.r.t. $\|\cdot\|_{1}$. Therefore $\tfrac{\lambda}{n}\sum_{j=1}^{n}h(p(x_{j}))$ is $\lambda$-strongly convex on the product simplex endowed with $\|\cdot\|_{\text{avg-}1}$. Since the surrogate usage term is convex (Assumption 1), $F_{S}$ is $\lambda$-strongly convex.
Lemma 4 (Uniform stability of ERM for $F_{S}$).
Under Assumptions 1 and 2, let $\hat{P}(S)$ and $\hat{P}(S^{(i)})$ be the ERM solutions. Then the solution mapping is uniformly stable:
$\big\|\hat{P}(S)-\hat{P}(S^{(i)})\big\|_{\text{avg-}1}~\leq~\mathcal{O}\!\Big(\frac{L+B}{\lambda\,n}\Big).$
Theorem 2 (Generalization bound for RSL-trained router).
Under Assumptions 1 and 2, the expected generalization gap satisfies
$\big|\;\mathbb{E}_{x}[\ell(\hat{P}(S);x)]~-~\tfrac{1}{n}\sum_{j=1}^{n}\ell(\hat{P}(S);x_{j})\;\big|~=~\mathcal{O}\!\Big(\frac{\rho\,(L+B)}{\lambda\,n}\Big),$
and, with probability at least $1-\delta$,
$\big|\;\mathbb{E}_{x}[\ell(\hat{P}(S);x)]~-~\tfrac{1}{n}\sum_{j=1}^{n}\ell(\hat{P}(S);x_{j})\;\big|~\leq~\mathcal{O}\!\Big(\frac{\rho\,(L+B)}{\lambda\,n}\Big)~+~\mathcal{O}\!\Big(\sqrt{\tfrac{\log(1/\delta)}{n}}\Big).$
Lemma 3 gives $\lambda$-strong convexity of $F_{S}$ under $\|\cdot\|_{\text{avg-}1}$. By a standard argument for strongly convex ERM with bounded differences, replacing one sample perturbs the objective by at most $O((L+B)/n)$ and yields uniform stability $\beta=\mathcal{O}((L+B)/(\lambda n))$ for the solution map (Lemma 4). The Lipschitz evaluation loss then converts stability into generalization: $\mathbb{E}[\text{gap}]\leq\rho\,\beta$, and a standard McDiarmid/Hoeffding argument implies the high-probability term $\mathcal{O}(\sqrt{\log(1/\delta)/n})$.
This matches the empirical robustness of RSL in low-data regimes: entropy regularization both improves optimization conditioning and reduces the hypothesis sensitivity to single-sample perturbations.
A.3 Additional experimental results.
The impact of Top-$K$. To explore the impact of Top-K routing, we conducted experiments on SST-2 and CoLA using Falcon-Mamba. The results are shown in Figure  5 . As $K$ increases from 1 to 3, we observe improved accuracy, suggesting that multiple experts can gain complementary information. However, further increasing $K$ may actually degrade performance. Therefore, how to set or dynamically learn the optimal $K$ value is a direction worthy of further research.
Figure 5: Top-K Routing Impact.
Cross-domain QA To evaluate LoRA-Mixer’s cross-domain generalization capabilities, we constructed two question-answering datasets: Medical-Mathematics and Mathematics-Coding. Each dataset contains 200 examples generated by DeepSeek-R1. These questions are more challenging. We conduct DeepSeek-R1 evaluations, and the results are shown in Table 10.
Table 10: Cross-domain performance of LoRA-Mixer on LLaMA3-8B.
Task Base LoRAHub MOLE MixLoRA LoRA-Mixer
Math-Medical 69.88 70.53 72.11 72.74 73.41
Math-Coding 59.37 61.08 62.24 63.10 63.46
A.4 Parameter analysis.
We tested the trainable parameters of LoRA-Mixer, MoLE, MixLoRA, and LoRAHub under LLaMA3-8B. The trainable parameters of LoRA-Mixer and MixLoRA are 3.88% and 8.08%, respectively (six experts + routing). LoRA-Mixer’s total parameters account for 48% of MixLoRA’s, with routing parameters accounting for 0.04%. MoLE’s training parameters are routing parameters. To ensure a fair comparison, we adjusted MoLE’s routing to reduce the trainable parameters to 0.04%. LoRAHub is a train-free method with 0% trainable parameters. PHATGOOSE’s trainable parameters are gating vectors, specifically 0.01%.
A.5 Datasets details.
We present information about all datasets used here for reference.
Table 11: Statistics of datasets used in experiments.
Dataset Train Dev Test Task
ARC-E 2251 570 2376 Multiple-choice QA
ARC-C 1119 299 1172 Multiple-choice QA
GSM8K 7473 1319 Math reasoning
MBPP 779 195 Code generation
CoLA 8551 1043 1063 Acceptability judgment
SST-2 10000 872 1821 Sentiment classification
Medical 10030 4004 4004 Medical QA
MRPC 3668 408 1725 Paraphrase detection
MNLI 392702 9815 9832 Natural language inference
QNLI 104743 5463 5463 Question answering/NLI
RTE 2490 277 3000 Natural language inference
QQP 363846 40430 390965 Paraphrase identification
PIQA 16113 1838 3277 Physical commonsense reasoning
HellaSwag 39905 10042 10042 Commonsense reasoning
BoolQ 9427 3270 3270 Boolean question answering
A.6 Computational Resources and Reproducibility Guarantee.
Our experiments were conducted on a Linux workstation equipped with a single NVIDIA A800 80GB GPU and a 32-core Intel Xeon CPU, using PyTorch version 2.1.0 and Transformers version 4.36.2. For the first stage of training, we used LoRA with parameters r = 64, alpha = 128, dropout = 0.1, batch size = 1, gradient accumulation = 4, and lr = 1e-5. The task loss function was cross-entropy loss, and a cosine scheduler with warmup ratio (ratio = 0.1) was used. In stage 2, RSL was added with $\alpha=0.1$ and $\lambda=0.5$ in RSL. $\beta$ of $\mathcal{L}_{\text{preserve}}$ was fixed to 5e-5, and a top-3 routing strategy was used. In the inference phase, we also adopt the top-3 routing strategy and test it in the same device environment.
A.7 Efficiency and Resource Overhead Analysis.
We tested inference/training time, and memory usage, and the results are shown in the table 12. These metrics highlight the advantages of LoRA-Mixer beyond task performance. Inference was performed using a single A800 80G GPU, with max_length=512 and max_new_tokens=15, to test single-sample inference time. To eliminate the effects of cold starts, all measurements were performed after a warm-up run.
Table 12: Inference time comparison across different methods.
Model Time (s)
LLaMA3-8B (Baseline) 0.441
LoRAHub 0.482
MoLE 0.563
MixLoRA 0.597
LoRA-Mixer (Ours) 0.574
LoRAHub uses a training-free approach, resulting in faster inference speed. MoLE is nearly identical to LoRA-Mixer. However, this coarse-grained approach often leads to poor performance. Overall, LoRA-Mixer achieves a better balance between inference speed and performance. Overall, LoRA-Mixer optimized with RSL achieves faster convergence and more stable training. For training information, LoRA-Mixer’s training is split into two stages: Stage 1 (expert training) and Stage 2 (router training). The results are shown in the table 13.
Table 13: Training stages with dataset size, time.
Stage Data Time (hours)
Stage 1 40k 8.86
Stage 2 2k 1.74
Stage 2 4k 3.38
LoRA-Mixer can skip stage 1 and quickly complete the stage 2 of training using very little data when computing resources are limited, and provide strong performance. This is very important in resource-constrained scenarios. For memory usage, we measure memory usage for 6 experts, r=64 (A800-80G, float32, batchsize=1, gradientaccumulation=4),the results are shown in the table 14:
Table 14: Memory usage comparison across different methods.
Method MixLoRA MoLE LoRA-Mixer
Memory Usage (GB) 43.07 36.57 38.48
A.8 Hyperparameter grid search.
The hyperparameters include $\alpha$$\beta$ and $\lambda$ are not many, and we design a simple yet effective strategy for tuning the hyperparameters. We start with heuristic values informed by prior work on multi-objective optimization: $\alpha=0.1$ and $\lambda=0.05$ in $\mathcal{L}_{\text{RSL}}$. We fix $\beta=5e-5$ in $\mathcal{L}_{\text{preserve}}$. These values are chosen to ensure that the task loss dominates the training in the early stages of training. We perform grid search over a narrow range of $\alpha$ and $\lambda$ values on a validation split of all training data (10% of the total data).
Table 15: Ablation study on $\alpha$ and $\lambda$. Best results are highlighted in bold.
$\alpha$ $\lambda$ Performance Load Balance (Variance) Routing Entropy
0.05 0.01 76.2% 0.12 1.9
0.1 0.05 78.5% 0.08 1.5
0.2 0.1 75.8% 0.05 1.2
The values $\alpha=0.1$ and $\lambda=0.05$ achieve the best trade-off across all metrics. To prevent RSL from interfering too much with task loss in the early stages of training, we dynamically adjust parameters during training, for example, gradually increasing $\alpha$ and $\lambda$ to their final values in the first 30% of epochs and then keeping them constant. This "warm-up" strategy stabilizes training and improves convergence.
A.9 Additional results on GLUE tasks.
We present the results of LoRA-Mixer on the QNLI and MNLI tasks, as shown in Table 16. The results show that LoRA-Mixer achieves state-of-the-art results on all GLUE tasks, demonstrating the powerful combination of RSL and LoRA-Mixer. The experimental settings are LLaMA3-8B and $r=16$, and the average performance over three runs is reported.
Table 16: Results on QNLI and MNLI. Higher is better.
Method QNLI MNLI
Base 90.36 89.95
LoRAHub 90.44 90.56
MOLE 91.48 91.04
MixLoRA 91.73 91.27
LoRA 91.34 90.87
LoRA-Mixer (Ours) 92.04 91.45
We exclude STSB because it focuses on measuring semantic similarity, which is inconsistent with our core goal of "cross-task expert collaboration" (classification, reasoning, and generation), and the results are more likely to be random. Furthermore, we want to be consistent with previous studies such as GOAT, LoRA-LEGO, and MixLoRA, which did not conduct experiments on this dataset.
A.10 Additional Empirical Support for LoRA Migration.
To better understand LoRA-Mixer’s cross-model transfer experiments, we analyzed the architectures of Mistral-7B and LLaMA3-8B. The Mistral-7B and LLaMA3-8B models share a core decoder-specific Transformer structure, and the key dimensions match, as shown below:
Table 17: Comparison of architecture components between Mistral-7B and LLaMA3-8B.
Architecture components Mistral-7B LLaMA3-8B
Number of layers 32 32
Hidden dimension ($d_{\text{model}}$) 4096 4096
Attention heads 32 32
Feedforward dimension ($d_{\text{ff}}$) 14336 14336
Attention projection layer Q, K, V, output (all $4096\rightarrow 4096$) Q, K, V, output (all $4096\rightarrow 4096$)
Activation function SwiGLU SwiGLU
This structural alignment ensures that linear projection layers with the same input/output dimensions can achieve parameter transfer.
A.11 Falcon-Mamba Architecture Analysis.
$h^{\prime}(t)$ $=\mathbf{A}h(t)+\mathbf{B}x(t),$ (16)
$y(t)$ $=\mathbf{C}h(t).$
$\mathbf{\overline{A}}$ $=\exp\left(\mathbf{\Delta A}\right),$ (17)
The discrete state-space equation with a step size of $\Delta$ becomes:
$h_{t}$ $=\mathbf{\overline{A}}h_{t-1}+\mathbf{\overline{B}}x_{t},$ (18)
$y_{t}$ $=\mathbf{C}h_{t}.$
$y$ $=x\otimes\mathbf{\overline{K}}.$
$x\xrightarrow{x\text{-proj}}(\Delta,B,C)$
$\Delta\xrightarrow{dt\text{-proj}}\Delta_{discretized}$
[…]
A.12 The impact of $r$ in LoRAs
@@
As shown in Table LABEL:r16,Table 19 and Table  20, within a certain range, as $r$ increases, the performance of the model can be improved to a certain extent. Our method is not only better than the basic model, but even better than the fine-tuned basic model on most tasks. This shows that our method can effectively combine existing knowledge through dynamic expert combination to form a more "intelligent" model.
A.13 Balance loss visualization.
@@
A.14 In-distribution experimental results.
Here, we present experimental results for LoRA-Mixer and PHATGOOSE on five in-distribution experiments (GSM8K, ARC-E, ARC-C, CoLA, and SST2). The results are shown in Table 21. Both methods use the same LoRA adapter (Q/K/V/O) on LLaMA3-8B and are trained using the same data; their routing strategies differ: LoRA-Mixer uses RSL, while PHATGOOSE uses sigmoid gating based on cosine similarity. It can be seen that LoRA-Mixer outperforms PHATGOOSE across the board.
Table 21: Comparison of Base, Phatgoose, and LoRA-Mixer across five datasets.
Dataset Base Phatgoose LoRA-Mixer
SST2 93.12 91.35 94.83
CoLA 79.14 78.27 81.17
GSM8K 57.92 58.87 62.66
ARC-E 88.45 87.13 89.47
ARC-C 78.65 75.94 79.89
In order to verify the enhanced performances of individual expert after LoRA-Mixer optimization. We selected four tasks, GSM8K, SST2, CoLA and HumanEval, for experiments. The results are shown in Table LABEL:tab:expert_ablation_gain_with_lora_reordered. We can find that LoRAs optimized by LoRA-mixer exhibit improved performance, especially in the GSM8K task, where the performance improved by 5.36% after adding mathematical experts. This result confirms the enhanced ability of each individual LoRA expert after LoRA-Mixer optimization.
A.15 Details of LoRA modules of Flan-T5.
To better understand the Flan-T5 LoRAs experiment from the Internet, we introduce the corresponding LoRAs configuration information here, details as 22:
Table 22: LoRA configuration details used in our experiments.